A child of mass 30 kg slides from rest down a wooden pole by a distance of 3.0 m. Calculate the average force of friction acting on the child if her speed is 2 m/s.
when she hits the ground.
Take g = 10m/s
1
Expert's answer
2021-08-25T18:13:20-0400

start by calculating the acceleration using the third equation of motion

"v^2=u^2+2as"

from rest u=0

"2^2=0^2+2(a)(3)"

"4=6a, a=0.67m\/s^2"

now "a=\frac{F_{net}}{m}=\frac{mg-friction force}{m}"

"ma=mg-friction force"

"friction force=ma-mg=30\times 0.67-30\times 10=-279.9N"

so the average friction force which is acting in the opposite direction is 279.9N